sum of three consecutive integers is -14.Find the numbers?
n , n+1=? , n+2=?

Answer by Alan3354(21605)

sum of three consecutive integers is -14.Find the numbers?
n , n+1=? , n+2=?
----------------
n + n+1 + n+2 = -14
3n+3 = -14
3n = -17
There is no integer solution. The sum of 3 consecutive integers will always be a multiple of 3.